OpenStudy (anonymous):

Simplify: (9 – 8)2 × (24 – 6 3)

OpenStudy (theviper):

Did u mean \[(9-8)^2\times(24-6)^3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no the second part does not have a ^2

OpenStudy (theviper):

If u meant \[(9-8)^2(24-6^3)\] than ur ans is\[(1^2)(24-216)\] \[=1\times-192=-192\]

OpenStudy (anonymous):

thank you :)